My problem is m=7,(4,1) I tied to solve a similar one which was
m=8,(7,5) and here is what I did.
y-5=m(x-x1)
y-5=(8)(x-x1)
y-5=(8)(x-(7)
y-5=8(x+7)
y-5=8x-14
y=8x-9
however something went wrong because the answer was 8x-51 and I have no idea how to do this problem. Please help me.

Answer by Alan3354(69443)

My problem is m=7,(4,1) I tied to solve a similar one which was
m=8,(7,5) and here is what I did.
y-5=m(x-x1)
y-5=(8)(x-x1)
y-5=(8)(x-(7)
----------------- Why did you change to +7 ?
y-5 = 8(x-7)
y-5 = 8x-56
y = 8x-51
------------
y-5=8(x+7) This was -7 ?
y-5=8x-14 Where did you get 14 ?
y=8x-9